OpenStudy (anonymous):

ty''-(t+1)y'+y=t^2 y1=e^t y2=t+1 differential equations 2nd order equations with variable coefficients

OpenStudy (anonymous):

try some general quadratic form

OpenStudy (anonymous):

is this supposed to be a Cauchy euler equation?

OpenStudy (anonymous):

not too sure of the names though :)

OpenStudy (anonymous):

what are those \(y_1,y_2\)? and @rperez36 yes it appears to be so

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually no I'm still confused. Where did you get \(y_1\) from?

OpenStudy (anonymous):

i think using laplace

OpenStudy (anonymous):

they are givens with the problem

OpenStudy (anonymous):

\[y_1=e^t \] \[y_2=t+1\]

OpenStudy (anonymous):

laplace transforms seem good but we dont have initial conditions

OpenStudy (anonymous):

I'm guessing you want us first to verify those two solutions satisfy the homogeneous equation $$ty''-(t+1)y'+y=0$$For \(y_1=e^t\) we have \(y=y'=y''=e^t\) hence:$$te^t-(t+1)e^t+e^t=0\\te^t-te^t-e^t+e^t=0\\0=0$$

OpenStudy (anonymous):

they both are solutions i think..then the general form must be a linear combination of the two

OpenStudy (anonymous):

For \(y_2=t+1\) we have \(y=t+1,y'=1,y''=0\) hence:$$-(t+1)+t+1=0\\0=0$$so indeed \(y_1,y_2\) are two linearly independent solutions to our homogeneous equation

OpenStudy (anonymous):

In Problems 37 through 40, use variation of parameters to find a general solution to the differential equation given that the functions y1 and y2 are linearly independent solutions to the corresponding homogeneous equation for Remember to put the equation in standard form

OpenStudy (anonymous):

@him1618 indeed but I believe she wants a particular solution to the non-homogeneous equation

OpenStudy (anonymous):

for t>0

OpenStudy (anonymous):

weve got to have a set of initial conditions for a particular solution i believe?

OpenStudy (anonymous):

@him1618 eh sorta but you're right that we're not interested in a specific solution instead a general one

OpenStudy (anonymous):

i would recommend laplace with assumptions that y(0) and y'(0) are 0

OpenStudy (anonymous):

we haven't gone over laplace :/

OpenStudy (anonymous):

then i think mrs perez id recommend taking a standard quadratic form, ax^2 + bx + c and putting it in place of y, and then equating coefficients to determine a ,b, c

OpenStudy (anonymous):

the solution is \[C_1e^t+C_2(t+1)-t^2\]

OpenStudy (anonymous):

you can try what i said and then linearly combine that with the 2 solutions given

OpenStudy (anonymous):

you need to use variation of parameters to find the particular solution. The given solutions are the corresponding solutions (the solutions that substitute to equal zero); the method of variation of parameters will help find the nonhomogenous part (that equals t^2). Variation of parameters indicates that the solution is of some form \[y_{p}=u_1y_1+u_2y_2\]

OpenStudy (anonymous):

id forgotten that one

OpenStudy (anonymous):

that wouldnt account for the t^2

OpenStudy (anonymous):

the only notes I have from this chapter are about second order Cauchy euler equations and every example he worked in class would =0. Just substituting in variables into the auxillary equation am^2+(b-a)m+c=0

OpenStudy (anonymous):

yes it accounts for the t^2. In fact, the instructions typed above indicates the method to use

OpenStudy (anonymous):

In Problems 37 through 40, use variation of parameters to find a general solution to the differential equation given that the functions y1 and y2 are linearly independent solutions to the corresponding homogeneous equation for Remember to put the equation in standard form

OpenStudy (anonymous):

@him1618 it very clearly states to use variation of parameters not the Laplace transform nor the method of undetermined coefficients

OpenStudy (anonymous):

i dont know the question..my bad then

OpenStudy (anonymous):

Have you ever used variations of parameters @rperez36?

OpenStudy (anonymous):

yeah it was the chapter before this so ill review the problem and try to get the solution

OpenStudy (anonymous):

\[u_1'=\frac{ -y_{2}f(x) }{ W }; u_2'=\frac{ y_1f(x) }{ W}\] where W is the Wronskian of the y_1 and y_2

OpenStudy (anonymous):

isn't this what i need to do first??

OpenStudy (anonymous):

you need to find the wronskian first.

OpenStudy (anonymous):

what about putting it in standard form? i did find the wronskian i got -e^t

OpenStudy (anonymous):

with the given y's anyway i thought that we needed the y's from the general solution first. ill give and example

OpenStudy (anonymous):

i got something different for the Wronskian. The given y's solve the homogenous part of the equation.

OpenStudy (anonymous):

if you plug in y=e^t into the differential equation the result is 0 not t^2

OpenStudy (anonymous):

@druminjosh what did you find for the Wronskian?

OpenStudy (anonymous):

http://prntscr.com/1i2qi4 that is one of our examples from variation of parameters chapter. the solution y1&y2 are from the complimentary solution...i think that is what it is called anyway

OpenStudy (anonymous):

and we also used the Y_c cos and sin for the determinates

OpenStudy (anonymous):

we're doing exactly what's done in the example. y_1 and y_2 given here ARE from the complimentary solution @rperez36

OpenStudy (anonymous):

sorry its kinda hard to see

OpenStudy (anonymous):

those big determinant bars in your example are the Wronskian

OpenStudy (anonymous):

ok that's what i though and used them for determinate so for the equation @druminjosh gave before for -y1 and y2 we use the givens again so it would be like -(t+1)t^2/W??

OpenStudy (anonymous):

almost the equations i gave above are for u_1' and u_2'

OpenStudy (anonymous):

i named it trash file so i'll know to throw it away later, not becaue your problem is trash :)

OpenStudy (anonymous):

The wronskian is -te^t

OpenStudy (anonymous):

lol ok and i made a dumb mistake i put t in bottom right instead of 1

OpenStudy (anonymous):

no worries, i figured it was a silly mistake

OpenStudy (anonymous):

ok one solution at a time. The general solution we are trying to find is of the form: \[y=u_1y_1+u_2y_2\]

OpenStudy (anonymous):

\[u_1'=\frac{ -y_2f(x) }{ W }\]

OpenStudy (anonymous):

you'll need to use integration by parts to find u_1

OpenStudy (anonymous):